# Pyramid-Monozone Synergistic Grasping Policy in Dense Clutter

### v1

@@ V-B2 Comparision with Baseline Methods in High-clutter
DexNet 4.0 and GraspNet are considered state-of-the-art (SOTA) for learning-based 4-DOF and 6D grasping, respectively. Therefore, to demonstrate the effectiveness of PMSGP’s grasping capability, we compare it with these two challenging methods in this section. It is important to note that because of the equipment limitations, we just compare the parallel-jaw version of DexNet 4.0 and the planar version of GraspNet. Moreover, we conducted experiments on high-clutter scenes composed of 50 ragdolls, 50 snacks, 50 toys, and 50 household objects, respectively. The experimental results are shown in II, indicate that PMSGP achieves GSR of 98.4% (250/254) for ragdolls, 94.3% (250/265) for snacks, 86.8% (250/288) for toys, and 86.2% (250/290) for household objects, all of which surpass DexNet 4.0 and GraspNet. More importantly, as the difficulty of the objects being grasped increases, the gap between PMSGP and the baseline methods becomes more obvious. For example, when grasping toys and household objects, PMSGP’s GSR exceeds theirs by up to 20%, demonstrating the high reliability and stability of our method.
@@ V-C1 Effectiveness of Pyramid Sequencing Policy
To emphasize the importance of TVA in PSP, we decompose PSP into TVA and Cross-prompted CPS for separate validation. For TVA, we use a mid-clutter scenario consisting of 20 household objects because PMSGP without TVA can only grasp within a very limited area. Additionally, PMSGP without TVA uses the best quality pixel as the prompt for segmentation, with other aspects consistent with PMSGP. For CPS, we use an extreme-clutter scenario consisting of 100 snacks, as their complex appearance helps to fully validate the effectiveness of CPS. Here, PMSGP without CPS uses the first segmentation, with other aspects also consistent with PMSGP too. The experimental results are shown in III and IV. The GSR of PMSGP without TVA is only 74.6% (100/134), whereas the original PMSGP achieves 90% (100/111), a gap of 15.4%, indicating that selecting a reasonable grasping order to avoid occlusions is crucial. Similarly, the GSR of PMSGP without CPS is 79.1% (500/632), compared with the original PMSGP’s 90% (500/564), proving the effectiveness of CPS. In addition, We show some grasping performance of PMSGP on objects of different grasping difficulties at different complexity in Fig 6. It can be seen that PMSGP is always able to avoid most occlusions and focus on the topmost objects for grasping.
@@ V-C2 Effectiveness of Monozone Sampling Policy
In this section, we demonstrate the effectiveness of MSP under the highest grasping difficulty, specifically by using extreme-clutter scenes composed of 100 household objects. These objects exhibit the greatest variation in materials, shapes, and appearances compared to other objects that we used. Additionally, the version without MSP uses quality-based sampling, while other aspects remain consistent with PMSGP. The experimental results, as shown in V, indicate that the GSR of PMSGP without MSP is 75% (500/667), compared to 84.9% (500/589) for PMSGP, a difference of approximately 10%, illustrating the obvious advantage of MSP over quality-based sampling in high-difficulty scenarios.
V-D Failure Case Analysis
In the above experiments, we performed more than 7,000 grasp attempts and achieved a total of 5,800 successful grasps. More importantly, we test PMSGP’s capability in extreme-clutter scenarios involving 100 novel objects. To the best of our knowledge, we have not seen any work that can make it. However, PMSGP still encounters some failures. The first is object slippage, which we plan to address by incorporating suction. The second is that due to the error of the depth camera, some grasps will slightly deviate from the topmost object. This can be achieved by implementing a high-precision industrial depth camera. The final involves objects with similar depths that are tightly packed together. This can be solved by designing a deformable gripper that deforms under force to separate objects.
TABLE III: The impact of with or without TVA
@@ V-D Failure Case Analysis
TABLE IV: The impact of with or without CPS
@@
TABLE V: The impact of with or without MSP
@@ VI Conclusion
In this paper, we propose the Pyramid-Monozone Synergistic Grasping Policy that enables robots to effectively avoid most occlusions in dense clutter grasping. By conducting extensive experiments, we demonstrate the effectiveness and reliability of our method. Future work can be divided into two major parts. The first part can focus on addressing the issues highlighted in the Failure Case Analysis to enhance the method proposed in this paper. The second part can involve using this method as a baseline and extending it to human-robot interaction for specific object retrieval. For instance, safely grasping a specific object in a cluttered scene without interfering with other objects, and securely handing it over to a person should be a fascinating point.

### v2

@@ V-B2 Comparision with Baseline Methods in High-clutter
DexNet 4.0 and GraspNet are considered state-of-the-art for learning-based 4-DOF and 6-DOF grasping, respectively. Therefore, to demonstrate the effectiveness of PMSGP’s grasping capability, we compare it with these two methods. It is important to note that we just compare the parallel gripper version of DexNet 4.0 and the planar version of GraspNet. Moreover, we conducted experiments in high-clutter scenes composed of 50 ragdolls, 50 snacks, 50 toys, and 50 household goods, respectively. The experimental results are shown in Table II, indicating that PMSGP achieves GSR of 98.4% (250/254) for ragdolls, 94.3% (250/265) for snacks, 86.8% (250/288) for toys, and 86.2% (250/290) for household goods. All surpassed DexNet 4.0 and GraspNet. More importantly, as the difficulty in grasping increases, the gap between PMSGP and the baseline methods becomes more obvious. For example, when grasping toys and household goods, PMSGP’s GSR exceeds theirs by up to 20%, demonstrating the high reliability of our method. We also visualize these results in Fig. 9. It is obvious that the bar length of our method is shorter than that of other methods in each subfigure, and it varies little across the trials and has smaller errors.
@@ V-C1 Effectiveness of Pyramid Sequencing Policy
To emphasize the importance of TVA in PSP, we decompose PSP into TVA and Cross-prompted CPS for separate validation. For TVA, we use a mid-clutter scenario consisting of 20 household goods, because PMSGP without TVA can only grasp within a very limited area. Additionally, PMSGP without TVA uses the best quality pixel as the prompt for segmentation, with other aspects consistent with PMSGP. For CPS, we use an extreme-clutter scenario consisting of 100 snacks, as their complex appearance helps fully validate the effectiveness of CPS. Here, PMSGP without CPS uses the first segmentation, with other aspects also consistent with PMSGP. The experimental results are shown in Table III and IV. The GSR of PMSGP without TVA is only 74.6% (100/134), whereas the original PMSGP achieves 90% (100/111), showing a substantial gap of 15.4%. It is therefore deemed crucial to select the right order of grasping targets to avoid occlusions. Similarly, the GSR of PMSGP without CPS is 79.1% (500/632), compared with the original PMSGP’s 90% (500/564), proving the effectiveness of CPS. In addition, we show some grasping performance of PMSGP on objects of different grasping difficulties at different complexity in Fig. 10. It can be seen that PMSGP is always able to avoid most occlusions and target the topmost objects.
@@ V-C2 Effectiveness of Monozone Sampling Policy
In this section, we demonstrate the effectiveness of MSP under the highest grasping difficulty, specifically in extremely cluttered scenes composed of 100 household goods. These objects exhibit the greatest variation in materials, shapes, and appearances compared to other objects that we use. Additionally, the version without MSP uses quality-based sampling, while other aspects remain consistent with PMSGP. The experimental results, as shown in Table V, indicate that the GSR of PMSGP without MSP is 75% (500/667), compared to 84.9% (500/589) for PMSGP. Two cases differ by approximately 10%, illustrating the obvious advantage of MSP over quality-based sampling in high-difficulty scenarios.
TABLE III: Impact of with or without TVA
@@
TABLE IV: Impact of with or without CPS
@@
TABLE V: Impact of with or without MSP
@@
V-D Failure Case Analysis
In the above experiments, we performed more than 7,000 grasp attempts and achieved a total of 5,800 successful grasps. More importantly, we tested PMSGP’s capability in extreme-clutter scenarios involving 100 novel objects, and the GSR is stable between 85% to 89%. To the best of our knowledge, no previous work has demonstrated similar performance. However, PMSGP still encounters some failures. The first issue is slippage during grasp execution, which occurs due to the smooth surface of the object. To address this, we plan to use parallel jaw grippers with high-friction finger pads or wrap the fingers with textured tape. Furthermore, errors from the depth camera can cause some grasps to deviate slightly from the topmost object. This problem can be mitigated by using a high-precision industrial depth camera. While our method reduces most occlusions during grasping, in certain cases, the lower part of the topmost object may still be occluded by the one beneath it. To optimize this, we are going to use amodal instance segmentation [36] to predict the occluded parts of the object, thereby getting the complete mask of the object. Finally, when objects with similar depths are tightly packed together, they will be difficult to grasp. This challenge can be overcome by designing a deformable parallel-jaw gripper that adapts under force to separate the objects, like [37].
@@ VI Conclusion
In this paper, we proposed the Pyramid-Monozone Synergistic Grasping Policy (PMSGP), which first employed the Pyramid Sequencing Policy (PSP) to determine the optimal grasping sequence and then applied the Monozone Sampling Policy (MSP) to identify the best grasp candidate. PMSGP enabled the robot to effectively mitigate occlusions in densely cluttered environments, significantly reducing the difficulty of grasping and enhancing the grasping success rate. We conducted over 7,000 real-world grasping attempts on 300 novel objects across various levels of clutter, including mid-clutter (20 objects), high-clutter (50 objects), and extreme-clutter (100 objects), demonstrating that PMSGP significantly outperforms seven competitive grasping methods.
Future work can be divided into two major parts. The first part can focus on addressing the issues highlighted in the Failure Case Analysis to enhance the method proposed in this paper. The second part can involve using this method as a baseline and extending it to human-robot interaction for specific object retrieval. For instance, safely grasping a specific object in a cluttered scene without interfering with other objects, and securely handing it over to a person should be noteworthy.